# Revisiting Transformers with Insights from Image Filtering

### v1

@@ Abstract
The self-attention mechanism, a cornerstone of Transformer-based state-of-the-art deep learning architectures, is largely heuristic-driven and fundamentally challenging to interpret. Establishing a robust theoretical foundation to explain its remarkable success and limitations has therefore become an increasingly prominent focus in recent research. Some notable directions have explored understanding self-attention through the lens of image denoising and nonparametric regression. While promising, existing frameworks still lack a deeper mechanistic interpretation of various architectural components that enhance self-attention, both in its original formulation [1] and subsequent variants. In this work, we aim to advance this understanding by developing a unifying image processing framework, capable of explaining not only the self-attention computation itself but also the role of components such as positional encoding and residual connections, including numerous later variants. We also pinpoint potential distinctions between the two concepts building upon our framework, and make effort to close this gap. We introduce two independent architectural modifications within transformers. While our primary objective is interpretability, we empirically observe that image processing-inspired modifications can also lead to notably improved accuracy and robustness against data contamination and adversaries across language and vision tasks as well as better long sequence understanding.
Index Terms:
Transformer, image processing, self-attention, positional encoding, residual connection, interpretability.
I Introduction
Transformers [1] have achieved state of the art performance across a wide variety of tasks in machine learning [2, 3] and, in particular, within natural language processing [4, 5] and computer vision [6, 7]
@@ I Introduction
The critical significance of understanding and controlling how Transformer works makes it vital to develop approachable theoretical frameworks both to explain currently available model variants as well as shed light on potential improvements that could be gained through such proof-friendly frameworks. Consequently, recent research has attempted to study Transformers, especially the self-attention mechanism, through variety of theoretical viewpoints such as kernel smoothing [8], dynamical systems [9], control theory [10], measure theory [11], variational denoising [12], and recently, interacting particle systems [13] to name but a few. Diversity of such perspectives reduces the friction in engagement of researchers from various backgrounds in studying Transformers.
In a complementary vein, the broader research community has already recognized the remarkable versatility of classical image processing methods [14, 15] in designing modern neural architectures. A particularly promising direction emerges from recent advancements in denoising algorithms applied in diffusion models [16]. Modern denoising diffusion probabilistic models (DDPMs) leverage iterative denoising refinements to generate high-fidelity images, and their underlying noise-reduction principles bear strong resemblance to filtering techniques in image processing. Moreover, nonlocal image processing techniques [17, 18] have found renewed relevance in deep learning. The fundamental principles behind these algorithms, including self-similarity, adaptive weighting, and spatially aware feature selection, bear strong conceptual parallels to the design of attention mechanisms [1, 19]. As a result, they offer valuable inspiration for refining the expressivity and efficiency of Transformer-based architectures in particular.
@@
In this work, we develop another theoretical perspective for Transformers that supports the understanding of the self-attention mechanism as well as positional encoding and residual connections through the lens of image processing algorithms, which some of the related prior works were lacking [12, 20].
@@
    We uncover a concrete connection between self-attention computation and image filtering and how an attention mechanism with properly utilized positional information could serve to make the filter bilateral (Section III).
@@
    We provide a specific, but natural, interpretation for residual connections via boosting of iterative image filtering algorithms (particularly by increasing signal-to-noise ratio (Proposition III.1)).
@@
    We theoretically prove that, with a bilateral filter-based incorporation of positional information (concatenated absolute PE or relative PE), long sequence stability of self-attention can be enhanced (Theorem III.2), which aligns with the empirical success of relative positional encoding over their absolute additive counterparts [21] (Remark III.4).
@@
    We show how a more sophisticated boosting technique can help to design residual connections to alleviate salient information flow through deep transformers (Section IV), and offer notable robustness improvements (Theorem IV.1).
@@
The paper is ordered as follows: In Section II, we provide a brief overview of background knowledge necessary for the topic. In Section III, we present our image filtering framework for Transformers by establishing a theoretical link between its components and image processing algorithms. In Section IV, we propose a novel residual connection that is intended to improve robustness by introducing input fidelity to the output of each layer, and consequently enhance robustness. In Section V, we provide experimental and empirical support for our theoretical observations and claims.
II Preliminaries
II-A Notations and Basic Definitions
Throughout this paper, we use the following notations. Lowercase Latin letters (e.g., $a,b,c$) represent scalar values. Bold lowercase Latin letters (e.g., $\bm{a},\bm{b}$) denote vectors. Unless stated otherwise, the coordinates of a vector $\bm{a}\in\mathbb{R}^{n}$ are denoted by $\bm{a}_{i}=a_{i}$ for $i\in[n]$, where $[n]:=\{1,2,\dots,n\}$. Bold uppercase Latin letters (e.g., $\bm{A},\bm{B}$) are used for matrices. The rows of a matrix $\bm{A}$ are denoted by $\bm{A}_{i}=\bm{a}_{i}$.
Definition II.1 (Inequality up to a Constant Factor).
@@ II-A Notations and Basic Definitions
$f(n)\lesssim g(n)\iff f(n)=O(g(n)).$ (1)
Definition II.2 (Asymptotic Equivalence).
@@
$f(n)\asymp g(n)\iff c_{1}g(n)\leq f(n)\leq c_{2}g(n).$ (2)
II-B Self-Attention Mechanism
@@
$=\sum_{j=1}^{N}\mathrm{softmax}\left(\frac{\bm{q}_{i}^{\top}\bm{k}_{j}}{\sqrt{d}}\right)\bm{v}_{j}$ (3)
@@
where $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{d\times d}$ are the corresponding learnable projection matrices. Eqn. II-B can then be expressed in matrix form as:
$\bm{U}=\mathrm{softmax}\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d}}\right)\bm{V},$ (4)
@@
As can be seen in Eqn. II-B, self-attention itself is position agnostic (e.g. it cannot account for positions of words in a sentence). For simplicity, we focus on the positional encoding of [1]. Given word embeddings $\bm{E}=[\bm{e}_{1},\dots,e_{N}]^{\top}\in\mathbb{R}^{N\times d}$, [1] have proposed to create positional vectors $\bm{p}_{i}$ using the sinusoidal function:
$\begin{cases}\bm{p}_{i,2t}=\sin\left(i/T^{2t/d}\right)\\ \bm{p}_{i,2t+1}=\cos\left(i/T^{2t/d}\right),\end{cases}$ (5)
@@
$\bm{x}_{i}=\bm{e}_{i}+\bm{p}_{i},\quad\quad\forall i\in[N],$ (6)
which is then fed into the self-attention block as input as given by Eqn. II-B. While the heuristic-driven design in Eqn. 6 has proven successful in practice, we shall show how it might actually be an overshoot towards leveraging positional information in later sections.
II-C Data-dependent Image Filters
The resemblance of Eqn. II-B to well-known data-dependent image filtering algorithms is not difficult to observe intuitively. Before delving into formal details, we provide a brief overview of two commonly adopted classes of data-dependent image filters in this section.
@@ II-C Data-dependent Image Filters
$y_{i}=u_{i}+\eta_{i},\quad\quad\forall i\in[n],$ (7)
@@
$\mathcal{J}_{K}(u(\bm{p}_{i}))=\sum_{j=1}^{n}[\underbrace{y_{j}-u(\bm{p}_{i})}_{\text{noise }\bm{\eta}}]^{2}K(\bm{p}_{i},\bm{p}_{j},y_{i},y_{j}),$ (10)
@@
$\frac{\partial\mathcal{J}_{K}}{\partial u}\bigg{|}_{\hat{u}}=\sum_{j=1}^{n}-2[y_{j}-\hat{u}(\bm{p}_{i})]K(\bm{p}_{i},\bm{p}_{j},y_{i},y_{j})=0.$ (11)
@@
Bilateral filter.
One special class of data-dependent image filters represented by Eqn. 12 is the Bilateral Filter [22], in which the kernel function is taken to account for spatial and photometric distances separably as:
$K_{BF}(\bm{p}_{i},$ $\bm{p}_{j},y_{i},y_{j}):=$
$\exp\left(-\frac{\|\bm{p}_{i}-\bm{p}_{j}\|^{2}}{h_{p}^{2}}\right)\exp\left(-\frac{(y_{i}-y_{j})^{2}}{h_{y}^{2}}\right),$ (13)
@@
Nonlocal Means (NLM) filter.
Another important example is the NLM filter [17], which essentially generalizes the Bilateral filter by replacing the individual pixel-wise proximity in Eqn. II-C with a patch-wise distance and discarding the effect of spatial distance as:
$K_{NLM}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j}):=\exp\left(-\frac{\|\bm{y}_{i}-\bm{y}_{j}\|^{2}}{h_{y}^{2}}\right).$ (14)
Patches could be vectorized to keep the distance metric consistent. This can also be formally regarded as letting $h_{p}\to\infty$ in the definition of $K_{BF}(\cdot)$ with patches $\bm{y}_{i}$ and $\bm{y}_{j}$ as last arguments. Although the spatial distance is ignored in theory, a bounded region around the position $\bm{p}_{i}$ is considered in practice to make the algorithm tractable [15].
III A Unifying Image Processing Framework for Transformer
To better isolate the contribution of self-attention, we focus on a simplified variant of the Transformer architecture that omits the intermediate MLP blocks typically stacked after each attention layer. A discussion of MLP layers and their role in the architecture is deferred to Section VI-B.
@@ III A Unifying Image Processing Framework for Transformer
III-A Problem framing for Transformer
@@
$\bm{y}_{i}=\bm{u}_{i}+\bm{\eta}_{i},\quad\quad\forall i\in[N],$ (15)
@@
Theorem III.1 ($1$-layer Transformer).
Consider a 1-layer transformer with a single attention head with identity projections $\bm{W}_{M:M\in\{Q,K,V\}}=\bm{I}$. Then, output vectors of its self-attention mechanism, as given by Eqn. II-B, are weighted least squares estimates of clean patches $\bm{u}_{i}$, given noisy patches $\bm{e}_{i}$ as Eqn. 15 with the following similarity kernel function:
$K_{SA}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})=\exp\left(\frac{(\bm{y}_{i}+\bm{p}_{i})^{\top}(\bm{y}_{j}+\bm{p}_{j})}{\sqrt{d}}\right).$ (17)
Remark III.1.
In Theorem III.1, we allow ourselves to take $\bm{W}_{M:M\in\{Q,K,V\}}=\bm{I}$ for simplicity. The more general case $\bm{W}_{M:M\in\{Q,K\}}\neq\bm{I}$ with proper constrains, however, can be related to learning a Mahalanobis distance to exploit the underlying data patterns to determine the most appropriate distance metric, which has proven its worth in robust imaging, regression, and deep learning tasks [23, 24, 25].
To understand an $n$-layer Transformer with $n$ self-attention blocks and residual connections through the lens of image processing, we first define the concept of boosting. “Boosting” in the image processing literature [26, 27, 15, 28] refers to using residual regularizing techniques enhancing image details and the signal-to-noise ratio (SNR) of the denoised image as defined below.
Definition III.1.
@@
When boosting is applied to denoising iterations, it effectively serves to maintain the amount of salient information exceedingly larger while allowing relatively smaller amount of noise induction. In fact, accommodating salient information flow in deep transformer layers is indeed necessary to avoid token uniformity, where all tokens tend to become almost indistinguishable, diminishing the representational capacity of the model [29].
Without loss of generality, we assume token vectors eventually collapse to a zero vector in Proposition III.1 since we can subtract the limiting vector otherwise. In case of multiple limiting vectors, tokens are grouped by their limit points and the same logic is applied group-wise. This enables us to essentially quantify “salient information” merely by their norms. Following the technique of [28], we now show that residual connection is indeed similar to a particular SNR boosting operation in a more general setting.
Proposition III.1 (Residual Connections Boost SNR).
@@
$\frac{\sigma(\bm{y}+\hat{\bm{y}})}{\sigma(\bm{y})}\geq\frac{\sqrt{1+2\alpha\beta+\alpha^{2}}}{1+\gamma}>1.$ (18)
All omitted nontrivial proofs are deferred to Appendix VII. Below we shall make the following important remarks to understand the assumption and the result of Proposition III.1:
Remark III.2.
The assumption that there exist constants $\alpha,\beta,\gamma\in[0,1]$ with $\min(\alpha,\beta)>\gamma$ such that $\|\hat{\bm{u}}\|=\alpha\|\bm{u}\|$, $\cos(\bm{u},\hat{\bm{u}})\geq\beta$, and $\|\hat{\bm{\eta}}\|=\gamma\|\bm{\eta}\|$ is fundamental to image denoising since it basically implies that the denoising output suppresses noise more than it reduces clean signal. For an ideal denoiser, one should have $\alpha=\beta=1$ and $\gamma=0$.
Remark III.3.
Notice that while $\sigma(\bm{y}+\hat{\bm{y}})>\sigma(\bm{y})$, it is generally not true that $\sigma(\bm{y}+\hat{\bm{y}})\geq\sigma(\hat{\bm{y}})$. However, even if SNR $\sigma(\hat{\bm{y}})$ happens to be greater than $\sigma(\bm{y}+\hat{\bm{y}})$ (e.g. when a denoising step was almost perfect), it does not guarantee that the absolute amount of clean signal is well-preserved. In fact, by assumption, $\|\hat{\bm{u}}\|=\alpha\|\bm{u}\|$ with $\alpha<1$. Then recursively applying this $n$ times, we get $\|\tilde{\bm{u}}\|\approx\alpha^{n}\|\bm{u}\|\to 0$ as $n\to\infty$, where $\tilde{\bm{u}}$ denotes the denoising estimate at iteration $n$ without residual connection. In Section IV, we discuss ways to overcome this issue by modifying the residual connection approach.
@@
Corollary III.1 ($n$-layer Transformer).
Consider an $n$-layer Transformer with each layer formed as in Theorem III.1 without the MLP blocks. Then, the computation flow of this transformer is equivalent to $n$-step iterative denoising with kernel $K_{SA}(\cdot)$ with boosting effect at each iteration via residual connections.
III-B Distinctions of Self-Attention
Despite the simplicity of the statement of Corollary III.1, it uncovers the hidden denoising task that self-attention is implicitly trying to solve. Nonetheless, there are a few potential concerns associated with the definition of $K_{SA}(\cdot)$ that we shall highlight through the following proposition:
Proposition III.2.

### v2

@@ Abstract
The self-attention mechanism, a cornerstone of Transformer-based state-of-the-art deep learning architectures, is largely heuristic-driven and fundamentally challenging to interpret. Establishing a robust theoretical foundation to explain its remarkable success and limitations has therefore become an increasingly prominent focus in recent research. Some notable directions have explored understanding self-attention through the lens of image denoising and nonparametric regression. While promising, existing frameworks still lack a deeper mechanistic interpretation of various architectural components that enhance self-attention, both in its original formulation [vaswani2017attention] and subsequent variants. In this work, we aim to advance this understanding by developing a unifying image processing framework, capable of explaining not only the self-attention computation itself but also the role of components such as positional encoding and residual connections, including numerous later variants. We also pinpoint potential distinctions between the two concepts building upon our framework, and make effort to close this gap. We introduce two independent architectural modifications within transformers. While our primary objective is interpretability, we empirically observe that image processing-inspired modifications can also lead to notably improved accuracy and robustness against data contamination and adversaries across language and vision tasks as well as better long sequence understanding.
keywords:
Transformer, image processing, self-attention, positional encoding, residual connection.
{MSCcodes}
68T07, 94A08
1 Introduction
Transformers [vaswani2017attention] have achieved state of the art performance across a wide variety of tasks in machine learning [khan2022transformers, lin2022survey] and, in particular, within natural language processing [al2019character, raffel2020exploring] and computer vision [touvron2021training, radford2021learning]
@@ 1 Introduction
The critical significance of understanding and controlling how Transformer works makes it vital to develop approachable theoretical frameworks both to explain currently available model variants as well as shed light on potential improvements that could be gained through such proof-friendly frameworks. Consequently, recent research has attempted to study Transformers, especially the self-attention mechanism, through variety of theoretical viewpoints such as kernel smoothing [tsai2019transformer], dynamical systems [lu2019understanding], control theory [Nguyen2024PIDformer], measure theory [vuckovic2020mathematicaltheoryattention], variational denoising [nguyen2023mitigating], and recently, interacting particle systems and mean-field theory [geshkovski2024mathematicalperspectivetransformers, geshkovski2024dynamic, bordelon2024infinite, bruno2025emergence] to name but a few. Diversity of such perspectives reduces the friction in engagement of researchers from various backgrounds in studying Transformers.
In a complementary vein, the broader research community has already recognized the remarkable versatility of classical image processing methods [milanfar2010isdenoising, milanfar2013tour, romano2017little] in designing modern neural architectures. A particularly promising direction emerges from recent advancements in denoising algorithms applied in diffusion models [Ho2020denoisingdiff, peebles2022scalable]. Modern denoising diffusion probabilistic models leverage iterative denoising refinements to generate high-fidelity images, and their underlying noise-reduction principles bear strong resemblance to filtering techniques in image processing. Moreover, nonlocal image processing techniques [buades2005nlm, gilboa2007nonlocal] have found renewed relevance in deep learning. The fundamental principles behind these algorithms, including self-similarity, adaptive weighting, and spatially aware feature selection, bear strong conceptual parallels to the design of attention mechanisms [vaswani2017attention, velickovic2018graph]. As a result, they offer valuable inspiration for refining the expressivity and efficiency of Transformer-based architectures in particular.
@@
In this work, we develop another theoretical perspective for Transformers that supports the understanding of the self-attention mechanism as well as positional encoding and residual connections through the lens of image processing algorithms, which some of the related prior works were lacking [nguyen2023mitigating, abdullaev2025transformer].
@@
    We uncover a concrete connection between self-attention computation and image filtering and how an attention mechanism with properly utilized positional information could serve to make the filter bilateral (Section 3).
@@
    We provide a specific, but natural, interpretation for residual connections via boosting of iterative image filtering algorithms (particularly by increasing signal-to-noise ratio (Proposition 3.1)).
@@
    We theoretically prove that, with a bilateral filter-based incorporation of positional information (concatenated absolute PE or relative PE), long sequence stability of self-attention can be enhanced (Theorem 3.2), which aligns with the empirical success of relative positional encoding over their absolute additive counterparts [press2022train] (Remark 3.4).
@@
    We show how a more sophisticated boosting technique can help to design residual connections to alleviate salient information flow through deep transformers (Section 4), and offer notable robustness improvements (Theorem 4.1).
@@
The paper is ordered as follows: In Section 2, we provide a brief overview of background knowledge necessary for the topic. In Section 3, we present our image filtering framework for Transformers by establishing a theoretical link between its components and image processing algorithms. In Section 4, we propose a novel residual connection that is intended to improve robustness by introducing input fidelity to the output of each layer, and consequently enhance robustness. In Section 5, we provide experimental and empirical support for our theoretical observations and claims.
2 Preliminaries
2.1 Notations and Basic Definitions
Throughout this paper, we use the following notations. Lowercase Latin letters (e.g., $a,b,c$) represent scalar values. Bold lowercase Latin letters (e.g., $\bm{a},\bm{b}$) denote vectors. Unless stated otherwise, the coordinates of a vector $\bm{a}\in\mathbb{R}^{n}$ are denoted by $\bm{a}_{i}=a_{i}$ for $i\in[n]$, where $n]:=\{1,2,\dots,n\}$. Bold uppercase Latin letters (e.g., $\bm{A},\bm{B}$) are used for matrices. The rows of a matrix $\bm{A}$ are denoted by $\bm{A}_{i}=\bm{a}_{i}$.
Definition 2.1 (Inequality up to a Constant Factor).
@@ 2.1 Notations and Basic Definitions
(1) $f(n)\lesssim g(n)\iff f(n)=O(g(n)).$
Definition 2.2 (Asymptotic Equivalence).
@@
(2) $f(n)\asymp g(n)\iff c_{1}g(n)\leq f(n)\leq c_{2}g(n).$
2.2 Self-Attention Mechanism
@@
(3) $=\sum_{j=1}^{N}\mathrm{softmax}\left(\frac{\bm{q}_{i}^{\top}\bm{k}_{j}}{\sqrt{d}}\right)\bm{v}_{j}$
@@
where $\bm{W}_{Q},\bm{W}_{K},\bm{W}_{V}\in\mathbb{R}^{d\times d}$ are the corresponding learnable projection matrices. Eqn. 2.2 can then be expressed in matrix form as:
(4) $\bm{U}=\mathrm{softmax}\left(\frac{\bm{Q}\bm{K}^{\top}}{\sqrt{d}}\right)\bm{V},$
@@
As can be seen in Eqn. 2.2, self-attention itself is position agnostic (e.g. it cannot account for positions of words in a sentence). For simplicity, we focus on the positional encoding of [vaswani2017attention]. Given word embeddings $\bm{E}=[\bm{e}_{1},\dots,e_{N}]^{\top}\in\mathbb{R}^{N\times d}$, [vaswani2017attention] have proposed to create positional vectors $\bm{p}_{i}$ using the sinusoidal function:
(5) $\begin{cases}\bm{p}_{i,2t}=\sin\left(i/T^{2t/d}\right)\\ \bm{p}_{i,2t+1}=\cos\left(i/T^{2t/d}\right),\end{cases}$
@@
(6) $\bm{x}_{i}=\bm{e}_{i}+\bm{p}_{i},\quad\quad\forall i\in[N],$
which is then fed into the self-attention block as input as given by Eqn. 2.2. While the heuristic-driven design in Eqn. 6 has proven successful in practice, we shall show how it might actually be an overshoot towards leveraging positional information in later sections.
2.3 Data-dependent Image Filters
The resemblance of Eqn. 2.2 to well-known data-dependent image filtering algorithms is not difficult to observe intuitively. Before delving into formal details, we provide a brief overview of two commonly adopted classes of data-dependent image filters in this section.
@@ 2.3 Data-dependent Image Filters
(7) $y_{i}=u_{i}+\eta_{i},\quad\quad\forall i\in[n],$
@@
(10) $\mathcal{J}_{K}(u(\bm{p}_{i}))=\sum_{j=1}^{n}[\underbrace{y_{j}-u(\bm{p}_{i})}_{\text{noise }\bm{\eta}}]^{2}K(\bm{p}_{i},\bm{p}_{j},y_{i},y_{j}),$
@@
(11) $\frac{\partial\mathcal{J}_{K}}{\partial u}\bigg|_{\hat{u}}=\sum_{j=1}^{n}-2[y_{j}-\hat{u}(\bm{p}_{i})]K(\bm{p}_{i},\bm{p}_{j},y_{i},y_{j})=0.$
@@
Bilateral filter
One special class of data-dependent image filters represented by Eqn. 12 is the Bilateral Filter [tomasi1998bf], in which the kernel function is taken to account for spatial and photometric distances separably as:
(13) $K_{BF}(\bm{p}_{i},\bm{p}_{j},y_{i},y_{j}):=$ $\exp\left(-\frac{\|\bm{p}_{i}-\bm{p}_{j}\|^{2}}{h_{p}^{2}}\right)\exp\left(-\frac{(y_{i}-y_{j})^{2}}{h_{y}^{2}}\right),$
@@
Nonlocal Means (NLM) filter
Another important example is the NLM filter [buades2005nlm], which essentially generalizes the Bilateral filter by replacing the individual pixel-wise proximity in Eqn. 13 with a patch-wise distance and discarding the effect of spatial distance as:
(14) $K_{NLM}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j}):=\exp\left(-\frac{\|\bm{y}_{i}-\bm{y}_{j}\|^{2}}{h_{y}^{2}}\right).$
Patches could be vectorized to keep the distance metric consistent. This can also be formally regarded as letting $h_{p}\to\infty$ in the definition of $K_{BF}(\cdot)$ with patches $\bm{y}_{i}$ and $\bm{y}_{j}$ as last arguments. Although the spatial distance is ignored in theory, a bounded region around the position $\bm{p}_{i}$ is considered in practice to make the algorithm tractable [milanfar2013tour].
3 A Unifying Image Processing Framework for Transformer
To better isolate the contribution of self-attention, we focus on a simplified variant of the Transformer architecture that omits the intermediate MLP blocks typically stacked after each attention layer. A discussion of MLP layers and their role in the architecture is deferred to Section 6.2.
@@ 3 A Unifying Image Processing Framework for Transformer
3.1 Problem framing for Transformer
@@
(15) $\bm{y}_{i}=\bm{u}_{i}+\bm{\eta}_{i},\quad\quad\forall i\in[N],$
@@
Theorem 3.1 ($1$-layer Transformer).
Consider a 1-layer transformer with a single attention head with identity projections $\bm{W}_{M:M\in\{Q,K,V\}}=\bm{I}$. Then, output vectors of its self-attention mechanism, as given by Eqn. 2.2, are weighted least squares estimates of clean patches $\bm{u}_{i}$, given noisy patches $\bm{e}_{i}$ as Eqn. 15 with the following similarity kernel function:
(17) $K_{SA}(\bm{p}_{i},\bm{p}_{j},\bm{y}_{i},\bm{y}_{j})=\exp\left(\frac{(\bm{y}_{i}+\bm{p}_{i})^{\top}(\bm{y}_{j}+\bm{p}_{j})}{\sqrt{d}}\right).$
Remark 3.1.
In Theorem 3.1, we allow ourselves to take $\bm{W}_{M:M\in\{Q,K,V\}}=\bm{I}$ for simplicity. The more general case $\bm{W}_{M:M\in\{Q,K\}}\neq\bm{I}$ with proper constrains, however, can be related to learning a Mahalanobis distance to exploit the underlying data patterns to determine the most appropriate distance metric, which has proven its worth in robust imaging, regression, and deep learning tasks [takeda2007lark, spira2007beltrami, nielsen2024elliptical].
To understand an $n$-layer Transformer with $n$ self-attention blocks and residual connections through the lens of image processing, we first define the concept of boosting. “Boosting” in the image processing literature [buhlmann2003boosting, osher2005iterative, milanfar2013tour, romano2015boosting] refers to using residual regularizing techniques enhancing image details and the signal-to-noise ratio (SNR) of the denoised image as defined below.
Definition 3.1.
@@
When boosting is applied to denoising iterations, it effectively serves to maintain the amount of salient information exceedingly larger while allowing relatively smaller amount of noise induction. In fact, accommodating salient information flow in deep transformer layers is indeed necessary to avoid token uniformity, where all tokens tend to become almost indistinguishable, diminishing the representational capacity of the model [dong2021pureattention].
Without loss of generality, we assume token vectors eventually collapse to a zero vector in Proposition 3.1 since we can subtract the limiting vector otherwise. In case of multiple limiting vectors, tokens are grouped by their limit points and the same logic is applied group-wise. This enables us to essentially quantify “salient information” merely by their norms. Following the technique of [romano2015boosting], we now show that residual connection is indeed similar to a particular SNR boosting operation in a more general setting.
Proposition 3.1 (Residual Connections Boost SNR).
@@
(18) $\frac{\sigma(\bm{y}+\hat{\bm{y}})}{\sigma(\bm{y})}\geq\frac{\sqrt{1+2\alpha\beta+\alpha^{2}}}{1+\gamma}>1.$
We shall make the following important remarks to understand the assumption and the result of Proposition 3.1 before presenting its proof:
Remark 3.2.
The assumption that there exist constants $\alpha,\beta,\gamma\in[0,1]$ with $\min(\alpha,\beta)>\gamma$ such that $\|\hat{\bm{u}}\|=\alpha\|\bm{u}\|$, $\cos(\bm{u},\hat{\bm{u}})\geq\beta$, and $\|\hat{\bm{\eta}}\|=\gamma\|\bm{\eta}\|$ is fundamental to image denoising since it esentially implies that the denoising output suppresses noise more than it reduces clean signal. For an ideal denoiser, one should have $\alpha=\beta=1$ and $\gamma=0$.
Remark 3.3.
[…]
Proof of Proposition 3.1. Using Definition 3.1, we have
(19) $\geq\frac{\|\bm{u}\|^{2}+\|\hat{\bm{u}}\|^{2}+2\|\bm{u}\|\|\hat{\bm{u}}\|\cos(\bm{u}\mid\hat{\bm{u}})}{\|\bm{\eta}\|^{2}+\|\hat{\bm{\eta}}\|^{2}+2\|\bm{\eta}\|\|\hat{\bm{\eta}}\|}\frac{\|\bm{\eta}\|^{2}}{\|\bm{u}\|^{2}}$
(20) $\geq\frac{(1+\alpha^{2})\|\bm{u}\|^{2}+2\alpha\beta\|\bm{u}\|^{2}}{(1+\gamma)\|\bm{\eta}\|^{2}+2\gamma\|\bm{\eta}\|^{2}}\frac{\|\bm{\eta}\|^{2}}{\|\bm{u}\|^{2}}$
(21) $=\frac{1+2\alpha\beta+\alpha^{2}}{(1+\gamma)^{2}},$
$\frac{\sigma(\bm{y}+\hat{\bm{y}})}{\sigma(\bm{y})}$ $\geq\frac{\sqrt{1+2\alpha\beta+\alpha^{2}}}{1+\gamma}$
$>\frac{\sqrt{1+2\gamma+\gamma^{2}}}{1+\gamma}$
$=1$
as desired.
@@
Corollary 3.1 ($n$-layer Transformer).
Consider an $n$-layer Transformer with each layer formed as in Theorem 3.1 without the MLP blocks. Then, the computation flow of this transformer is equivalent to $n$-step iterative denoising with kernel $K_{SA}(\cdot)$ with boosting effect at each iteration via residual connections.
3.2 Distinctions of Self-Attention
Despite the simplicity of the statement of Corollary 3.1, it uncovers the hidden denoising task that self-attention is implicitly trying to solve. Nonetheless, there are a few potential concerns associated with the definition of $K_{SA}(\cdot)$ that we shall highlight through the following proposition:
Proposition 3.2.